OpenStudy (anonymous):

trinomial factoring of x^3y^3+8z^3

OpenStudy (anonymous):

like \(x^3+b^3=(a+b)(a^2-ab+b^2)\) so long as you recognize that \(x^3y^3=(xy)^3\) and also that \(8z^3=(2z)^3\) so you can use the above with \(a=xy, b=2z\)

OpenStudy (anonymous):

sorry i meant like \(a^3+b^3=(a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

what letters go with what numbers

OpenStudy (anonymous):

x^3y^3 goes with a^3

OpenStudy (anonymous):

thank you